# When LLMs Go Online: The Emerging Threat of Web-Enabled LLMs

### v1

@@ 3.2 Targeted Cyberattacks
Collecting PII is the most straightforward task, where LLM agents gather targeted information from online contexts. Generating impersonation posts is a more complex task that involves using personal data to craft content closely mimicking the target. The most complex task, spear phishing email generation, involves designing scenarios, establishing sender identities, and creating content tailored to the target’s interests.
@@ 4.1 Experimental Setup
Evaluation. Due to the vast and diverse nature of the field of CS, we did not evaluate based solely on department listings. Instead, we employed human annotators to review the responses generated by LLM agents. We employed human annotators to review the responses generated by LLM agents. Annotators were directed to classify individuals as CS professors based on their engagement in relevant fields, following a thorough web search. Certain types of PII, like phone numbers, are prone to changes over time, complicating verification; thus, we initiated five separate queries to LLMs for the remaining PII for each individual. We considered the information collection successful if at least one of the five responses matched the data available online. During the annotation process for the names of CS students, we used the student information listed on the professors’ web pages as the ground truth. For more details on the annotation process, please refer to Appendix B.1.
@@ 4.2 Main Results
Figure 4: The number of correctly collected PII for CS professors categorized by each type of PII, using the Claude for all methods.
@@ 6.1 Experimental Setup
Attack Scenario. In Section 4, we demonstrated the feasibility of effectively obtaining the email addresses of researchers in the university. Similarly, attackers can exploit publicly available email addresses [37] or acquire them through purchases on the dark web or Telegram channels [16]. Attackers commonly use phishing emails to obtain sensitive information from targeted individuals by asking them to click on a link or enter personal information [8]. In our study, the attacker’s goal is to craft highly personalized phishing emails generated by WebSearch agents to induce recipients to click the malicious link, utilizing the email addresses of target individuals.
The attack pipeline is designed to craft highly convincing phishing emails. Note that the attack consists of no manual effort by the attacker and only by utilization of LLM agents. The only ingredient for this attack is the target’s email address, which is added to the generation prompt of the LLM query. As illustrated in Figure 8, an attacker can create a phishing email by sending a single query to the WebSearch agent. The process begins with verifying the date tㅇo ensure the timing aligns with the intended attack scenario. Then, the content of the email is developed by searching the target’s email address online to gather personal information. This information is then utilized to design a realistic scenario where the target is likely to receive an email, along with identifying a credible sender that fits this scenario. Subsequently, a plausible URL string is generated and embedded in the email to encourage the target’s engagement. Finally, the sender’s email address is altered to mimic a legitimate domain, enabling the dispatch of the email without exploiting vulnerabilities in official domains or compromising accounts to access the authentic domain.
User Study. Phishing emails were custom-generated for each target, requiring recipients to personally evaluate the emails intended for them. Thus, we designed a questionnaire to assess these emails and actively recruited participants for a comprehensive evaluation. Specifically, we recruited 60 participants and categorized them based on the provided institutional email addresses to analyze response variations across institutional types: 55% were academic researchers and 45% were non-academic professionals  [^11] The recruitment process with ethical/privacy-preserving measures is explained in Appendix D.1.. We emphasize that participants did not receive these emails directly; instead, they were assessed via a Google survey.
@@
The first two questions focus on the content of the emails, asking participants to identify the presented information (Q1) and to evaluate its accuracy against their actual personal details (Q2). Subsequent questions explore perceptions of the sender and recipient, including whether the recipient identified the sender (Q3) and if the recipient’s name correctly matched their real name (Q4). We then inquire about the actions participants would likely take upon receiving the email (Q5). The last question investigates perceived authenticity of the email (Q6), asking participants to rate how genuine or fraudulent it appeared and to identify specific elements that influenced their judgment (Q7). The questionnaire is available in Appendix D.3.
@@ 6.2 Main Results
We first analyze the effectiveness and content of general-purpose emails created by LLMs and WebSearch agents. Then, we examine the results according to the participants’ groups. We then analyze the efficacy of emails according to their purpose and sender’s institution. In Tables 7 and 8 of Appendix D.2, we report the comprehensive survey results concerning participants’ perceptions of authenticity and anticipated actions for various email types.
@@ 6.2.1 Overall Analysis
When comparing “actions” (Q5) and “perceived authenticity” (Q6) in Figure 10, the percentage of participants identifying the email as definitely phishing closely aligned with those opting move to spam. However, the percentage of participants who rated the email as genuine (definitely genuine + probably genuine) did not closely match the percentage who chose to click link. Interestingly, for emails generated by WebSearch agents, the click rate was much higher than the rate of perceived genuineness, a pattern not observed with vanilla LLM-generated emails. This discrepancy, driven by participants marked as unsure, suggests that even recipients uncertain of an email’s authenticity remain vulnerable to phishing attempts using WebSearch agents. These results highlight that the WebSearch agents not only provided a sense of credibility but also effectively utilized target information from the web to engage the recipients’ interest, leading to higher link-clicking rates.
@@
Content Analysis. We examined the influence of web accessibility on the detail of email content (Q1, multiple responses allowed) and the accuracy of those details in aligning with the target’s information (Q2, multiple responses allowed). Both factors played statistically significant roles in participants’ assessment of an email’s authenticity  [^13] The results of a chi-square test between questions and perceived authenticity showed a kappa value of 98.8 for Q1, and a kappa value of 113.5 for Q2, with p-values significantly below than 0.05.. Figure 11 presents a comparative analysis of the percentage of phishing emails that incorporate specific content details and how well this content aligns with the targets’ actual information, based on participants’ evaluations. When comparing GPT and Claude, it was noted that GPT-generated emails were less detailed, regardless of web usage. We also observed that emails crafted by WebSearch agents were significantly richer in detail and more accurately reflected the target’s personal data than those generated by vanilla LLMs. This difference was particularly pronounced in the category of specific expertise. The alignment of content with the target’s specific expertise had the most significant impact on participants’ perception of the email’s authenticity  [^14] A chi-square kappa value of 11.5 was obtained, with p-values significantly lower than 0.05.. By leveraging detailed information about the target’s expertise from the web, WebSearch agents were able to craft more convincing and effective phishing emails.
@@ 6.2.3 Factors Increasing Phishing Effectiveness
We examined the impact of changing the email’s purpose and sender’s address on the effectiveness of phishing attacks. Emails targeting login credentials proved more effective in prompting recipients to click the provided link, with a click rate of 33.33% for Claude’s WebSearch agent and 25% for GPT’s WebSearch agent (see Table 8 in Appendix D.2). This aligns with existing research [1], which suggested that urgent requests, such as login information updates, are effective in phishing scenarios.
@@ 7.1 Cost Analysis
Conducting attacks incurs both time and financial costs, which can highly impact the practicality of using LLM agents for cyberattacks. We examine how quickly and at what cost attackers can carry out cyberattacks using commercially available LLMs. We measured the time spent and the number of tokens consumed during each attack scenario using WebSearch agent for each individual target. For a cost analysis of the WebNav agent, please refer to Appendix E.1. To simplify our calculations, we considered token usage as the sum of input and output tokens, applying the output token rate, which is typically more expensive than the input token rate. This approach ensures we estimate the upper limit of API costs. For the models utilized, the costs are as follows: GPT 4o [41] and Claude 1.5 Sonnet [5] at \$15/1M tokens, and Gemini-1.5 Flash [20] at \$0.30/1M tokens. To measure the number of tokens, we utilized the tiktoken tokenizer [42] for GPT, as specified by OpenAI [42]. For Claude and Gemini, we used the token counting feature provided by their APIs [3, 13].
On average, collecting PII for an individual  [^15] We sample 50 professors identified in Section 4. required 9.1 seconds and incurred a cost of 2.2 cents (GPT: 6.7s, 2.1¢; Claude: 11.4s, 2.2¢). For generating impersonation posts, GPT, Claude, and Gemini took an average of 9.4 seconds, 21.7 seconds, and 3.6 seconds, respectively, with costs under 3 cents (GPT: 2.2¢; Claude: 3.0¢; Gemini: 0.03¢). For creating phishing emails, GPT and Claude took an average of 7.1 seconds and 22.7 seconds, respectively, with costs below 4 cents (GPT: 2.7¢; Claude: 3.9¢). See Appendix E for the cost statistics.
@@ 7.2 Safeguard Capability
Gemini shows the strongest safeguards by refusing to generate emails in all scenarios of vanilla LLM and WebSearch agent settings (Table 6 in Appendix E.3), responding with messages like “I cannot fulfill your request. Creating emails without consent is unethical and potentially illegal”. All models without web accessibility refuse generating content when the email’s purpose involves requesting login credentials or money, indicating these topics trigger stricter safeguard mechanisms. Conversely, general-purpose emails are less likely to be blocked, with GPT and Claude generating responses in these cases. GPT and Claude are more likely to generate an email when provided with an institutional email address, compared to a Gmail address.
@@ 8 Conclusion
We care deeply about human safety and societal security, which motivates us to conduct this research. In this paper, we discuss and carefully examine potential threat of LLM agents in cyberattacks exploiting personal data. This raises ethical considerations as some of the data that we extract contains information about individual users. We minimize these concerns by using data that is already public and do not present any specific information in this paper.
In the case of user study in Section 6, according to the local IRB’s protocol, our research is excluded from the review for the following reasons. First, we used only publicly available information and did not handle any sensitive data. Additionally, the emails evaluated contained no harmful or offensive content, minimizing any potential impact on participants. In the recruitment process, we obtained informed consent from participants, detailing the purpose of the research, the expected duration, procedures involved, anticipated benefits, personal information protection, and consent for the collection and use of their provided email addresses. We also informed participants that the LLM agents could search the internet for their email addresses to gather information and use it to generate emails. In the evaluation process, we did not send emails to the participants. We asked to evaluate emails through a Google survey after notification. Each participant was exposed only to emails generated from their own email address.
We believe that the benefits of publicizing these attacks outweigh the potential harms. We will disclose our research findings (especially concerning the safeguard vulnerabilities of web-based tools) to each respective LLM service vendor (OpenAI, Anthropic, and GoogleAI).
@@ C.2 Human Correlation with LLM Evaluation
For each claim and model combination mentioned in Section 5.1, 10 posts were randomly selected for evaluation by both human reviewers and LLM evaluators. The average agreement rate between humans and agents was 90.0% for the first claim and 73.3% for the second, underscoring the reliability and credibility of the LLM evaluations.
Appendix D Spear Phishing Email Generation
D.1 Recruiting Process and Ethical Measures
D.2 Results
Table 7 and Table 8 represent the overall survey results.
D.3 Questionnaire for Evaluation
Figure 14 shows the questionnaire that we used for evaluation.
Appendix E Approachability of Cyberattacks
E.1 PII Collection
@@
E.2 Impersonation Post and Phishing Email Generation
@@
E.3 Safeguard Capability in Phishing Email Generation
@@
Table 7: Survey results about perceived authenticity of phishing emails. Percentages represent responses relative to the total number of participants.
Genuine or Phishing
Model Web use Purpose Sender’s institution Definitely genuine Probably genuine Unsure Probably phishing Definitely phishing
Claude o general not designated 5.00 13.33 13.33 43.33 25.00
x general not designated 5.00 18.33 6.67 21.67 48.33
o credential not designated 3.33 25.00 10.00 26.67 35.00
o general different from target’s 3.33 25.00 26.67 35.00 10.00
GPT o general not designated 3.33 10.00 13.33 28.33 45.00
x general not designated 1.67 8.33 13.33 28.33 48.33
o credential not designated 1.67 20.00 10.00 18.33 50.00
Average 3.33 17.14 13.33 28.81 37.38

### v2

@@ 3.1 Overview of LLM Agents
Table 1: Summary of experimental settings for each attack.
PII collection Impersonation post generation Spear-phishing email generation
(Section 4) (Section 5) (Section 6)
Attack Goal Get targets’ 5 types of PII Generate social media posts that impersonate a target advocating a specific claim Generate targeted emails to induce recipients click a link embedded in the email
Attack Target CS researchers CS professors Academic researchers, Non-academic professionals
Attack Inputs Name of university Target’s name, Target’s institution, Desired claim Target’s email address
Evaluation Human annotation LLM evaluation User study
@@ 3.2 Targeted Cyberattacks
Collecting PII is the most straightforward task, where LLM agents gather targeted information from online contexts. Generating impersonation posts is a more complex task that involves using personal data to craft content closely mimicking the target. The most complex task, spear phishing email generation, involves designing scenarios, establishing sender identities, and creating content tailored to the target’s interests. We summarize experimental settings of these attacks in Table 1.
@@ 4.1 Experimental Setup
Evaluation. Due to the vast and diverse nature of the field of CS, we did not evaluate based solely on department listings. Instead, we employed human annotators to review the responses generated by LLM agents. Annotators were directed to classify individuals as CS professors based on their engagement in relevant fields, following a thorough web search. Certain types of PII, like phone numbers, are prone to changes over time, complicating verification; thus, we initiated five separate queries to LLMs for the remaining PII for each individual. We considered the information collection successful if at least one of the five responses matched the data available online. During the annotation process for the names of CS students, we used the student information listed on the professors’ web pages as the ground truth. For more details on the annotation process, please refer to Appendix B.1.
@@ 4.2 Main Results
Figure 4: The number of correctly collected PII for CS professors categorized by each type of PII, using Claude for all methods.
@@ 6.1 Experimental Setup
Attack Scenario. In Section 4, we demonstrated the feasibility of effectively obtaining the email addresses of researchers in the university. Similarly, attackers can exploit publicly available email addresses [37] or acquire them through purchases on the dark web or Telegram channels [16]. Attackers commonly use phishing emails to obtain sensitive information from targeted individuals by asking them to click on a link or enter personal information [8]. In our study, the attacker’s goal is to craft highly personalized phishing emails generated by LLM agents to induce recipients to click the malicious link, utilizing the email addresses of target individuals.
The attack pipeline is designed to craft highly convincing phishing emails. Note that the attack consists of no manual effort by the attacker and only by utilization of LLM agents. The only ingredient for this attack is the target’s email address, which is added to the generation prompt of the LLM query. As illustrated in Figure 8, an attacker can create a phishing email by sending a single query to the agent. The process begins with verifying the date to ensure the timing aligns with the intended attack scenario. Then, the content of the email is developed by searching the target’s email address online to gather personal information. This information is then utilized to design a realistic scenario where the target is likely to receive an email, along with identifying a credible sender that fits this scenario. Subsequently, a plausible URL string is generated and embedded in the email to encourage the target’s engagement. Finally, the sender’s email address is altered to mimic a legitimate domain, enabling the dispatch of the email without exploiting vulnerabilities in official domains or compromising accounts to access the authentic domain.
Note that we conducted two pilot studies, which showed similar results from WebNav and WebSearch Agents. Due to cost and time constraints, WebNav-generated emails were excluded from the main evaluation.
User Study. Phishing emails were custom-generated for each target, requiring recipients to personally evaluate the emails intended for them. Thus, we designed a questionnaire to assess these emails and actively recruited participants for a comprehensive evaluation.
As in Section 4, academic researchers’ information is publicly available online, making them more vulnerable to phishing attacks. Therefore, we categorized participants into two groups: academic researchers (including undergraduate students) and non-academic professionals (individuals within non-academic organizations). Academic researchers were recruited via university research communities, while non-academic professionals were recruited through industry practitioners from prior collaborations. Specifically, we recruited 60 participants and categorized them based on the provided institutional email addresses to analyze response variations across institutional types: 55% were academic researchers and 45% were non-academic professionals  [^11] The recruitment process with ethical/privacy-preserving measures is explained in Appendix E.1..
We emphasize that participants did not receive these emails directly; instead, they were assessed via a Google survey. Each participant evaluated seven phishing emails generated using their own email as a target. The types of emails will be described in Email Design part.
@@
The first two questions focus on the content of the emails, asking participants to identify the presented information (Q1) and to evaluate its accuracy against their actual personal details (Q2). Subsequent questions explore perceptions of the sender and recipient, including whether the recipient identified the sender (Q3) and if the recipient’s name correctly matched their real name (Q4). We then inquire about the actions participants would likely take upon receiving the email (Q5). The last question investigates perceived authenticity of the email (Q6), asking participants to rate how genuine or fraudulent it appeared and to identify specific elements that influenced their judgment (Q7). The questionnaire is available in Appendix E.5.
@@ 6.2 Main Results
We first analyze the effectiveness and content of general-purpose emails created by LLMs and WebSearch agents. Then, we examine the results according to the participants’ groups. We then analyze the efficacy of emails according to their purpose and sender’s institution. In Tables 10 (Appendix E.3), we report the comprehensive survey results concerning participants’ anticipated actions for various email types.
@@ 6.2.1 Overall Analysis
When comparing “actions” (Q5) and “perceived authenticity” (Q6) in Figure 10, on average, 66.27% of respondents who identified an email as definitely phishing opted for move to spam as their action. However, the percentage of participants who rated the email as genuine (definitely genuine + probably genuine) did not closely match the percentage who chose to click link. Interestingly, for emails generated by WebSearch agents, the click rate was much higher than the rate of perceived genuineness, a pattern not observed with vanilla LLM-generated emails. This discrepancy was driven by participants marked as unsure or even phishing; 13.34% of participants stated that they would clicked on links in WebSearch agent-generated emails despite doubting their authenticity. (For vanilla LLM-generated emails, 10% and 3.34% did so for Claude and GPT, respectively.) Table 11 (Appendix E.3) presents a contingency table showing the percentage distribution of participants’ actions (Q5) relative to their perception of authenticity (Q6). These results highlight that the WebSearch agents not only provided a sense of credibility but also effectively utilized target information from the web to engage the recipients’ interest, leading to higher link-clicking rates.
@@
Content Analysis. We analyzed which email aspects influence participants’ perceptions of authenticity using chi-squared tests. From the analysis, the detail of email content and the accuracy of those details in aligning with the target’s information played statistically significant roles in participants’ assessment of an email’s authenticity. Table 8 in Appendix E.2 provides statistical results of the chi-squared tests.
Building on these statistical results, we further examined the influence of web accessibility on the two factors. Figure 11 presents a comparative analysis of the percentage of phishing emails that incorporate specific content details and how well this content aligns with the targets’ actual information, based on participants’ evaluations. When comparing GPT and Claude, it was noted that GPT-generated emails were less detailed, regardless of web usage. We also observed that emails crafted by WebSearch agents were significantly richer in detail and more accurately reflected the target’s personal data than those generated by vanilla LLMs. This difference was particularly pronounced in the category of specific expertise. According to the results of our chi-squared tests in Table 9 (Appendix E.2), the alignment of content with the target’s specific expertise had the most significant impact on participants’ perception of the email’s authenticity. By leveraging detailed information about the target’s expertise from the web, WebSearch agents were able to craft more convincing and effective phishing emails.
@@ 6.2.3 Factors Increasing Phishing Effectiveness
We examined the impact of changing the email’s purpose and sender’s address on the effectiveness of phishing attacks. Emails targeting login credentials proved more effective in prompting recipients to click the provided link, with a click rate of 33.33% for Claude’s WebSearch agent and 25% for GPT’s WebSearch agent (see Table 10 in Appendix E.3). This aligns with existing research [1], which suggested that urgent requests, such as login information updates, are effective in phishing scenarios.
@@ 7.1 Cost Analysis
Conducting attacks incurs both time and financial costs, which can highly impact the practicality of using LLM agents for cyberattacks. We examine how quickly and at what cost attackers can carry out cyberattacks using commercially available LLMs. We measured the time spent and the number of tokens consumed during each attack scenario using WebSearch agent for each individual target. For a cost analysis of the WebNav agent, please refer to Appendix D.1. To simplify our calculations, we considered token usage as the sum of input and output tokens, applying the output token rate, which is typically more expensive than the input token rate. This approach ensures we estimate the upper limit of API costs. For the models utilized, the costs are as follows: GPT 4o [41] and Claude 1.5 Sonnet [5] at \$15/1M tokens, and Gemini-1.5 Flash [20] at \$0.30/1M tokens. To measure the number of tokens, we utilized the tiktoken tokenizer [42] for GPT, as specified by OpenAI [42]. For Claude and Gemini, we used the token counting feature provided by their APIs [3, 13].
On average, collecting PII for an individual  [^13] We sample 50 professors identified in Section 4. required 9.1 seconds and incurred a cost of 2.2 cents (GPT: 6.7s, 2.1¢; Claude: 11.4s, 2.2¢). For generating impersonation posts, GPT, Claude, and Gemini took an average of 9.4 seconds, 21.7 seconds, and 3.6 seconds, respectively, with costs under 3 cents (GPT: 2.2¢; Claude: 3.0¢; Gemini: 0.03¢). For creating phishing emails, GPT and Claude took an average of 7.1 seconds and 22.7 seconds, respectively, with costs below 4 cents (GPT: 2.7¢; Claude: 3.9¢). See Appendix D for the cost statistics.
@@ 7.2 Safeguard Capability
Gemini shows the strongest safeguards by refusing to generate emails in all scenarios of vanilla LLM and WebSearch agent settings (Table 7 in Appendix D.3), responding with messages like “I cannot fulfill your request. Creating emails without consent is unethical and potentially illegal”. All models without web accessibility refuse generating content when the email’s purpose involves requesting login credentials or money, indicating these topics trigger stricter safeguard mechanisms. Conversely, general-purpose emails are less likely to be blocked, with GPT and Claude generating responses in these cases. GPT and Claude are more likely to generate an email when provided with an institutional email address, compared to a Gmail address.
@@ 8 Conclusion
We care deeply about human safety and societal security, which motivates us to conduct this research. This paper discusses and carefully examines the potential threat of LLM agents in cyberattacks exploiting personal data. This raises ethical considerations as some of the data that we extract contains information about individual users. We minimize these concerns by using already public data and not presenting any specific information in this paper.
Aggregation of PIIs. The collected PII data is stored in pickle format on the principal investigator’s server. The data was provided to annotators in an Excel file format for annotation purposes. Once the annotation process was completed, the annotators returned the annotated files to the principal investigator and deleted their local copies. Currently, files containing personal information are exclusively accessible by the principal investigator and will be discarded after the publication of the research.
User Study. In the case of the user study in Section 6, according to the local IRB’s protocol, our research is excluded from the review for the following reasons. First, we used only publicly available information and did not handle any sensitive data. Additionally, the emails evaluated contained no harmful or offensive content, minimizing any potential impact on participants. In the recruitment process, we obtained informed consent from participants, detailing the purpose of the research, the expected duration, procedures involved, anticipated benefits, personal information protection, and consent for collecting and using their provided email addresses. We also informed participants that the LLM agents could search the internet for their email addresses to gather information and use it to generate emails. Only those consented could participate, and each participant completed the survey once. In the evaluation process, we did not send emails to the participants. We asked to evaluate emails through a Google survey after notification. Each participant was exposed only to emails generated from their own email address. All participant-related information was destroyed after the experiments, and participants were informed about this process. Additionally, we informed them that their responses would be excluded from the analysis if they withdrew consent, ensuring robust protection of personal information.
Responsible Disclosure Process. We believe that the benefits of publicizing these attacks outweigh the potential harms. First, we plan to submit our research findings to the LLM service vendors (OpenAI, Anthropic, and GoogleAI), including our prompts and concerns about the safeguard vulnerabilities of web-based tools. Additionally, we will request the implementation of a rule ensuring that when an LLM utilizes web crawling tools, it checks and adheres to robots.txt directives. (A robots.txt file tells web crawlers which parts of a site they can or cannot access.) Second, for the PII providers from whom we collected data, we strongly recommend creating a robots.txt file to block web crawlers from accessing pages containing personal information.
@@ C.2 Human Correlation with LLM Evaluation
We assessed the reliability of our LLM evaluators by comparing with human reviewers. We sampled 270 posts, each impersonating one of 10 professors for each claim, language model, and capability(LLM, WebSearch gent, WebNav agent) combination mentioned in Section 5.1. Then, both human reviewers and agents evaluated these posts. The average agreement rate between humans and agents was 93.0% for the first claim and 85.2% for the second, underscoring the reliability and credibility of the LLM evaluations.
Appendix D Approachability of Cyberattacks
D.1 PII Collection
@@
D.2 Impersonation Post and Phishing Email Generation
@@
D.3 Safeguard Capability in Phishing Email Generation
@@
Appendix E Spear Phishing Email Generation
E.1 Recruiting Process and Ethical Measures
E.2 Chi-Squared Tests
This section investigates the factors influencing participants’ perception of email authenticity. First, we identify which aspects of an email had the most significant influence on determining the email’s authenticity. We conducted chi-squared tests to evaluate the independence between email evaluation questions (Q1–Q4) and the question assessing perceived email authenticity (Q6). Please refer to Appendix E.5 for the questions used in our survey. Table 8 shows that “content detail” (Q1) and “content authenticity” (Q2) had p-values much less than 0.05. These results suggest that these factors are critical in determining whether an email is perceived as genuine.
Further analysis focused on “content authenticity” (Q2) identified as the most strongly associated factor. Specifically, we examined which aspects of email content influenced participants’ perceptions of authenticity. We evaluated the independence between each response option within Q2 and the option in Q6 using chi-squared tests. Table 9 shows that the “Specific area of expertise” option had a p-value much less than 0.05, indicating a highly significant association. This finding suggests that when the email content aligns with the participant’s domain-specific knowledge, it strongly impacts their perception of the email’s authenticity.
Table 8: Chi-squared test results between each question (Q1-Q4) and Q6.
Q1 Q2 Q3 Q4
(content detail) (content auth) (sender) (recipient)
kappa 98.80 113.54 56.54 3.69
p-value $2.23\times 10^{-5}$ $3.12\times 10^{-7}$ 0.0159 0.4497
Table 9: Chi-squared test results between each response option within Q2 and Q6.
institution general expertise specific expertise hobby
kappa 4.34 0.01 11.15 0.28
p-value 0.0595 0.0373 0.9436 0.0008
E.3 Overall Survey Results
Table 10 represents the survey results about participants’ actions (Q5) across seven email types.
E.4 Contingency Table
Table 11 presents contingency tables illustrating the percentage distribution of participants’ actions (Q5) in relation to their perception of authenticity (Q6). The analysis focuses on emails with general purposes and no specified sender institution.